# Tag-LLM: Repurposing General-Purpose LLMs for Specialized Domains

### v1

@@ 1 Introduction
We evaluate Tag-LLM on a diverse set of ten domains, encompassing eight languages and two specialized scientific domains (protein sequences and SMILES molecule representations). We chose the latter two because they represent non-linguistic, highly specialized domains with distributions significantly different from natural language, and they are relatively well understood with established baselines. We experiment with the LLaMA-7B model (Touvron et al., 2023). Our results demonstrate that the domain tags can act as effective context switchers, and a function tag can be applied to multiple domains to solve different tasks, achieving zero-shot generalization to unseen problems. We show that Tag-LLM improves LLaMA’s performance on ten translation tasks, matching specialized multilingual LLMs. We also achieve state-of-the-art results on multiple drug discovery datasets from the Therapeutics Data Commons benchmark (Huang et al., 2021). Given its model-agnostic and plug-and-play nature, Tag-LLM provides a systematic way to leverage the rapid advancements in LLMs to tackle a wider range of real-world problems. The code is available at Tag-LLM.
@@ 2.3 LLMs for Specialized Non-Linguistic Domains
Extensive efforts have been devoted to developing domain-specific models for specialized tasks such as coding (Ahmad et al., 2020), mathematical reasoning (Zong & Krishnamachari, 2023), protein structure prediction (Jumper et al., 2021), and gene effect prediction (Tu et al., 2022; Shen et al., 2022). Yet these models typically have custom architectures and are trained from scratch, requiring large amounts of in-domain data and compute. Here we seek a method that can leverage existing LLMs and requires minimal in-domain data, allowing for faster on-demand adaptation by non-domain-experts. In a similar vein, several recent works adapt pretrained LLMs to non-NLP fields such as physical and biological sciences (Singhal et al., 2022; Vinod et al., 2023; Shen et al., 2023), tabular data (Dinh et al., 2022), and quantitative reasoning (Lewkowycz et al., 2022). These methods require either fine-tuning the entire model (Lu et al., 2022a; Shen et al., 2023) or manually converting non-text data into text (Dinh et al., 2022). In contrast, we learn continuous embeddings that condition the model with domain-specific information, which is more parameter-efficient than fine-tuning and more effective than text conversion (see experimental results in Section 5).
@@ 4 Specialized NLP: Multilingual Translation
Seen Language Pairs | Seen Domains / Unseen Pair | Unseen Domains | |||||||||
en-fr | fr-en | en-ru | ru-en | en-de | de-en | it-fr | fr-it | es-pt | pt-es | ||
General | ours | 40.5 | 45.7 | 14.2 | 30.1 | 23.0 | 40.5 | 28.8 | 23.9 | 24.6 | 27.8 |
Pretrained | gpt-3 | 36.1 | 42.8 | 11.2 | 28.1 | 25.9 | 40.4 | - | - | - | - |
Specialized | xglm | 36.0 | 40.4 | 24.2 | 30.4 | 27.6 | 38.8 | - | - | - | - |
Pretrained | bloom | 45.0 | 45.6 | - | - | - | - | 31.4 | 24.0 | 29.1 | 28.1 |
Supervised | m2m | 42.0 | 37.2 | 27.1 | 27.5 | 32.6 | 35.8 | 34.4 | 28.6 | 28.1 | 26.9 |
@@
# Trainable | Single-Domain, Single-Instance | Single-Domain, Multi-Instance | Multi-Domain | |||
Params | Descriptor | QED | Drug Combination | Binding Affinity | ||
MSE () | MSE () | MAE () | Pearson () | |||
Tuning-Based | ours | 86K | 0.005 | 0.008 | 12.21 | 0.527 |
lora | 1M | 0.006 | 0.020 | 13.46 | 0.125 | |
prompt tuning | 86K | 0.011 | 1.387 | 17.65 | 0.054 | |
linear probing | 4K | 0.041 | 0.012 | 24.11 | 0.180 | |
Tuning-Free | hard prompt | 0 | 0.236 | 0.298 | 23.39 | 0.017 |
Specialized | nearest neighbor | 0 | 0.012 | 0.040 | - | - |
@@
plm-based | 4M | - | - | - | 0.538 | |
mlp-based | 13M | - | - | 16.85 | 0.433 |
@@ 5.1 Single-Domain, Single-Instance Tasks
We start with two single-domain, single-instance tasks: descriptor prediction for proteins, and Quantative Estimation of Drug-likeness (QED) score prediction for chemical compounds. The ground truth labels are automatically generated using the Python peptides and RDKit libraries, respectively. Since the tasks involve numerical prediction, we train the function tags and the paired regression heads using the MSE loss. We compare with three sets of baselines: (1) PEFT methods, including LoRA (Hu et al., 2021), prompt tuning (Lester et al., 2021), and linear probing; (2) hard-prompting with task instructions; and (3) a non-parametric nearest-neighbor algorithm that looks for similar training data points and leverages their labels to make predictions. For more details about the tasks and baselines, see Appendix A.2.
@@ 5.3 Multi-Domain, Multi-Instance Tasks
ours | 86K | 0.005 | 0.008 | 12.21 | 0.527 | |
lora | + domain in text | 1M | 0.008 | 0.015 | 12.79 | 0.552 |
prompt tuning | + domain in text | 86K | 0.007 | 0.0085 | 14.73 | 0.37 |
linear probing | + domain in text | 4K | 0.049 | 0.011 | 14.64 | 0.31 |
lora | + reg head | 1M | 0.007 | 0.015 | 12.53 | 0.534 |
prompt tuning | + reg head | 86K | 0.008 | 0.0083 | 15.29 | 0.38 |
linear probing | + reg head | 4K | 0.041 | 0.012 | 24.11 | 0.18 |

### v3

@@ 1 Introduction
We evaluate Tag-LLM on a diverse set of ten domains, encompassing eight languages and two specialized scientific domains (protein sequences and SMILES molecule representations). We chose the latter two because they represent non-linguistic, highly specialized domains with distributions significantly different from natural language, and they are relatively well understood with established baselines. We experiment with the LLaMA-7B model (Touvron et al., 2023). Our results demonstrate that the domain tags can act as effective context switchers, and a function tag can be applied to multiple domains to solve different tasks, achieving zero-shot generalization to unseen problems. We show that Tag-LLM improves LLaMA’s performance on ten translation tasks, matching specialized multilingual LLMs. We also achieve state-of-the-art results on multiple drug discovery datasets from the Therapeutics Data Commons benchmark (Huang et al., 2021). Given its model-agnostic and plug-and-play nature, Tag-LLM provides a systematic way to leverage the rapid advancements in LLMs to tackle a wider range of real-world problems. Code is available at https://github.com/sjunhongshen/Tag-LLM.
@@ 2.3 LLMs for Specialized Non-Linguistic Domains
Extensive efforts have been devoted to developing domain-specific models for specialized tasks such as coding (Ahmad et al., 2020), mathematical reasoning (Zong & Krishnamachari, 2023), protein structure prediction (Jumper et al., 2021), and gene effect prediction (Tu et al., 2022; Shen et al., 2022). Yet these models typically have custom architectures and are trained from scratch, requiring large amounts of in-domain data and compute. Here we seek a method that can leverage existing LLMs and requires minimal in-domain data, allowing for faster on-demand adaptation by non-domain-experts. In a similar vein, several recent works adapt pretrained LLMs to non-NLP fields such as physical and biological sciences (Singhal et al., 2022; Vinod et al., 2023; Shen et al., 2023, 2024), tabular data (Dinh et al., 2022; Roberts et al., 2021), and quantitative reasoning (Lewkowycz et al., 2022). These methods require either fine-tuning the entire model (Lu et al., 2022a; Shen et al., 2023) or manually converting non-text data into text (Dinh et al., 2022). In contrast, we learn continuous embeddings that condition the model with domain-specific information, which is more parameter-efficient than fine-tuning and more effective than text conversion (see experimental results in Section 5).
@@ 4 Specialized NLP: Multilingual Translation
Seen Language Pairs | Seen Domains / Unseen Pair | Unseen Domains | ||||||||
en-fr | fr-en | en-ru | ru-en | en-de | de-en | it-fr | fr-it | es-pt | pt-es | |
Specialized | ||||||||||
M2M | 42.0 | 37.2 | 27.1 | 27.5 | 32.6 | 35.8 | 34.4 | 28.6 | 28.1 | 26.9 |
XGLM | 36.0 | 40.4 | 24.2 | 30.4 | 27.6 | 38.8 | - | - | - | - |
BLOOM | 45.0 | 45.6 | - | - | - | - | 31.4 | 24.0 | 29.1 | 28.1 |
General-Purpose | ||||||||||
GPT-3 | 36.1 | 42.8 | 11.2 | 28.1 | 25.9 | 40.4 | - | - | - | - |
Tag-LLaMA (Ours) | 40.5 | 45.7 | 14.2 | 30.1 | 23.0 | 40.5 | 28.8 | 23.9 | 24.6 | 27.8 |
@@
Single-Domain, Single-Instance | Single-Domain, Multi-Instance | Multi-Domain | ||||
Fine-Tuning | # Trainable | Descriptor | QED | Drug Combination | Binding Affinity | |
Model | Method | Params | MSE () | MSE () | MAE () | Pearson () |
Specialized | ||||||
Nearest Neighbor | 0 | 0.012 | 0.040 | - | - | |
@@
PLM | 4M | - | - | - | 0.538 | |
MLP | 13M | - | - | 16.85 | 0.433 | |
Direct Prompting | ||||||
LLaMA | 0 | 0.236 | 0.298 | 23.39 | 0.017 | |
GPT-4 | 0 | 0.031 | 0.18 | 22.35 | -0.1051 | |
Galactica | 0 | 0.089 | 0.36 | 22.4 | -0.0053 | |
Text+Chem T5 | 0 | - | 0.31 | 26.60 | - | |
LlaSMol | 0 | - | 0.061 | 24.15 | - | |
Fine-Tuning Based | ||||||
Galactica | Linear Probing | 2K | 0.022 | 0.13 | 22.76 | 0.1557 |
Text+Chem T5 | Linear Probing | 768 | - | 0.048 | 22.22 | - |
LlaSMol | Linear Probing | 4K | - | 0.016 | 21.43 | - |
LLaMA | Linear Probing | 4K | 0.041 | 0.012 | 24.11 | 0.180 |
LLaMA | Prompt Tuning | 86K | 0.011 | 1.387 | 17.65 | 0.054 |
LLaMA | LoRA | 1M | 0.006 | 0.020 | 13.46 | 0.125 |
Tag-LLaMA (Ours) | Tag-LLM | 86K | 0.005 | 0.008 | 12.21 | 0.527 |
@@ 5.1 Single-Domain, Single-Instance Tasks
We start with two single-domain, single-instance tasks: descriptor prediction for proteins, and Quantative Estimation of Drug-likeness (QED) score prediction for chemical compounds. The ground truth labels are automatically generated using the Python peptides and RDKit libraries, respectively. Since the tasks involve numerical prediction, we train the function tags and the paired regression heads using the MSE loss. We compare with three sets of baselines. First, we directly prompt state-of-the-art LLMs with task instructions. We consider five models: the general-purpose GPT-4 (OpenAI, 2023) and LLaMA-7B, the science-focused Galactica-1.3B (Taylor et al., 2022), the chemistry-specialized LlaSMol (Yu et al., 2024) and Text+Chem T5-Base (Christofidellis et al., 2023). Second, we apply PEFT methods including LoRA (Hu et al., 2021), prompt tuning (Lester et al., 2021), and linear probing to LLaMA-7B. We also apply linear probing to the specialized LLMs mentioned earlier. Lastly, we consider task-specific, non-LLM models and non-parametric methods, such as the nearest-neighbor algorithm that looks for similar training data points and leverages their labels to make predictions. For more details about the tasks and baselines, see Appendix A.2.
@@
Tag-LLM | 86K | 0.005 | 0.008 | 12.21 | 0.527 | |
LoRA | + Text-Form Domain Info | 1M | 0.008 | 0.015 | 12.79 | 0.552 |
Prompt Tuning | + Text-Form Domain Info | 86K | 0.007 | 0.0085 | 14.73 | 0.37 |
Linear Probing | + Text-Form Domain Info | 4K | 0.049 | 0.011 | 14.64 | 0.31 |
LoRA | + Regression Head | 1M | 0.007 | 0.015 | 12.53 | 0.534 |
Prompt Tuning | + Regression Head | 86K | 0.008 | 0.0083 | 15.29 | 0.38 |
Linear Probing | + Regression Head | 4K | 0.041 | 0.012 | 24.11 | 0.18 |
@@ 6.2 Ablations on Regression Head
In all of the above experiments, we use LLaMA-7B as the LLM backbone to showcase the efficacy of our method. In Appendix A.3, we further investigate applying Tag-LLM to another open-source model, Mistral-7B (Jiang et al., 2023). The results demonstrate the broad applicability of Tag-LLM beyond just the LLaMA models.
@@ 7 Conclusion and Future Work
Impact Statement
This paper calls for ML community’s attention to take advantage of LLMs and apply them to a wider range of real-world problems beyond the NLP domains. This moves towards truly democratizing machine learning in real life. In terms of broader societal impact, our work can exert a positive influence as it contributes to reusing existing models and resources, reducing the computational burden of developing new large-scale models on massive data. However, lowering the barrier for applying LLMs to a wide range of tasks necessarily comes with the risk of misuse. Hence, it is imperative to develop adaptation methods with better privacy, safety, and fairness guarantees.
@@ A.2.3 Generation setup
A.3 Tag-Mistral
For the results reported in the main text, we mainly used LLaMA-7B as the LLM backbone to showcase the efficacy of our method. Here, we further investigate applying Tag-LLM to another open-source model, Mistral-7B (Jiang et al., 2023). The following results demonstrate the broad applicability of Tag-LLM beyond just the LLaMA models.
Descriptor (MSE, ) | QED (MSE, ) | DC (MAE, ) | BA (Pearson , ) | |
Tag-LLaMA | 0.005 | 0.008 | 12.21 | 0.527 |
Tag-Mistral | 0.0048 | 0.011 | 12.34 | 0.472 |
Best Task-Specific Expert Model | 0.012 | 0.04 | 16.85 | 0.588 |
Across all four tasks, augmenting Mistral with input tags outperforms the PEFT baselines and the domain-specific LLM baselines in Table 2, showing the effect of 1) injecting domain knowledge into the LLM via domain tags and 2) leveraging additional supervised data via function tags. Meanwhile, both Tag-Mistral and Tag-LLaMA perform comparably well. This demonstrates the compatibility of Tag-LLM with diverse pretrained backbones.